# BeamLLM: Vision-Empowered mmWave Beam Prediction with Large Language Models

### v1

@@ Abstract
In this paper, we propose BeamLLM, a vision-aided millimeter-wave (mmWave) beam prediction framework leveraging large language models (LLMs) to address the challenges of high training overhead and latency in mmWave communication systems. By combining computer vision (CV) with LLMs’ cross-modal reasoning capabilities, the framework extracts user equipment (UE) positional features from RGB images and aligns visual-temporal features with LLMs’ semantic space through reprogramming techniques. Evaluated on a realistic vehicle-to-infrastructure (V2I) scenario, the proposed method achieves $61.01\%$ top-$1$ accuracy and $97.39\%$ top-$3$ accuracy in standard prediction tasks, significantly outperforming traditional deep learning models. In few-shot prediction scenarios, the performance degradation is limited to $12.56\%$ (top-$1$) and $5.55\%$ (top-$3$) from time sample $1$ to $10$, demonstrating superior prediction capability.
@@ I Introduction
Millimeter-wave (mmWave) communication has garnered significant attention due to its abundant spectrum resources above $26$ GHz, enabling high-speed data transmission. However, the high operating frequency results in substantial path loss. To address this challenge, massive multiple-input multiple-output (mMIMO) antenna arrays are extensively employed, which utilize highly directional beamforming techniques to mitigate propagation losses. Furthermore, the short wavelength of mmWave signals facilitates compact antenna spacing, which enables the integration of large-scale antenna arrays within constrained physical dimensions. The effectiveness of directional beamforming depends on precise alignment between transmit and receive beams. Beam training addresses this challenge by scanning predefined codebooks at both the transmitter and receiver to identify the optimal beam pair, thereby maximizing received signal power without requiring the exhaustive acquisition of full channel state information (CSI).
Compared to legacy sub-6 GHz MIMO systems, beam training in mmWave systems faces heightened challenges: 1) Large antenna arrays lead to high-dimensional channel matrices, increasing training overhead; 2) Frequent beam tracking, especially in high-mobility scenarios (e.g., vehicle-to-everything (V2X) and unmanned aerial vehicles (UAVs)), introduces prohibitive latency. Recent studies [1, 2, 3, 4] have explored sensing-aided beam prediction, leveraging multimodal sensor data such as RGB images, radar, LiDAR, and GPS to improve efficiency and reduce training overhead. As a key enabler of integrated sensing and communication (ISAC) in 6G, this approach holds significant potential to enhance the performance of mmWave MIMO systems.
@@
To maintain beam prediction performance, deep learning (DL) is commonly used to extract user equipment (UE) movement features from received sensing data, enabling more accurate future beam selection. Due to its powerful non-linear feature extraction capability, DL has been widely explored in wireless communication tasks, including channel estimation[5] and beam prediction. Recent breakthroughs in large language models (LLMs), such as GPT-4[6] and DeepSeek[7], have demonstrated remarkable contextual reasoning and few-shot generalization abilities. While LLMs are originally designed for natural language processing (NLP), LLMs have shown strong cross-modal learning capabilities, thus extending their applications to time series for forecasting and computer vision (CV) tasks.
@@
Inspired by these advantages of the LLMs, several methods applying LLMs have been proposed for channel prediction[8], beam prediction [9], and port prediction for fluid antennas[10]. Built on these developments, in this paper, we propose a vision-aided beam prediction framework, named BeamLLM, which utilizes LLMs to process RGB images, thereby enabling more efficient and adaptive beam selection. Unlike [9], our method does not rely on historical beam indices or angle of departure (AoD) information. Instead, BeamLLM relies solely on visual features for beam prediction. Furthermore, to ensure robust performance and practical applicability, we validate our framework using real-world measurement datasets, to demonstrate its potential for deployment in real-world scenarios.
@@
The rest of this paper is organized as follows: Section II provides a system model and problem formulation of the beam prediction task. The proposed BeamLLM framework is presented in Section III. Section IV presents extensive simulation results, including performance comparisons with benchmark methods, along with detailed discussions. Finally, we conclude our work in Section V.
Figure 1: Illustration of the system model considered.
@@ II-A System Model
Fig. 1 illustrates the system model considered for vehicle-to-infrastructure (V2I) mmWave communication. In this model, the base station (BS) deploys a mmWave phased-array receiver with $M$ elements of half-wavelength spacing and an RGB camera. The antenna array enables the BS to perform beamforming, while the camera captures images within its field of view at a certain frame rate for sensing and downstream applications.
@@
We assume that the BS has a predefined beamforming codebook =$\mathcal{F}=\{\mathbf{f}_{1},\cdots,\mathbf{f}_{|\mathcal{F}|}\}$, containing $|\mathcal{F}|$ beams, where $\mathbf{f}_{m}\in\mathbb{C}^{N\times 1},m=1,\cdots,M$ represents the $m$-th beamforming vector. The UE is assumed to have a single antenna. At time step $t$, the user transmits a single symbol $s[t]\in\mathbb{C}$ that satisfies the power constraint $\mathbb{E}[|s[t]|^{2}]=P$, where $P$ represents the transmit power. At the BS, the received signal $y[t]$ can be expressed as:
$y[t]=\mathbf{h}^{H}[t]\mathbf{f}_{m}[t]s[t]+n[t],$ (1)
where $\mathbf{h}[t]$ is the channel vector, $\mathbf{f}_{m}[t]$ is the $m$-th beamforming vector from the codebook in time step $t$, and $n[t]\sim\mathcal{CN}(0,\sigma^{2})$ is the additive white Gaussian noise (AWGN) with variance $\sigma^{2}$.
@@ II-B Problem Formulation
This paper mainly focuses on the beam prediction problem at the BS. Given the available sensing information up to time $t-1$, the BS attempts to determine the optimal beams for $H\in\mathbb{Z}^{+}$ future time steps, specifically for $t,\cdots,(t+H-1)$. We define the optimal beam at time step $t$ as the beam that provides the highest beamforming gain, given by:
$\mathbf{f}_{m}^{*}[t]=\arg\max_{\mathbf{f}_{m}[t]\in\mathcal{F}}|\mathbf{h}^{H}[t]\mathbf{f}_{m}[t]|^{2}.$ (2)
When perfect CSI knowledge is unavailable, beam training serves as an alternative method for determining the optimal beam. However, with a narrow beam codebook, the training overhead can be significant, and the likelihood of identifying the optimal beam is often low when the pre-beamforming SNR is poor. Because the optimal beam selection at the transmitter and receiver depends on the surrounding environment of the transceiver, our work aims to leverage visual information at the BS to assist beam selection and develop a beam prediction framework.
@@ III Large Laguage Model-Based Beam Prediction
In this section, we introduce BeamLLM to tackle the vision-assisted beam prediction task outlined in Section II. Fig. 2 illustrates the proposed BeamLLM. The architecture mainly comprises three components, i.e., the visual data feature extraction module and the backbone module.
@@ III-A Visual Data Feature Extraction Module
To process raw RGB data for the vision-aided beam prediction task, we employ the YOLOv4 object detector[11]. This detector identifies potential UEs within RGB images and extracts bounding box vectors $\mathbf{b}$. For a single image $\mathbf{X}_{I}\in\mathbb{R}^{W\times H\times C}$, bounding box vector is given as:
$\mathbf{b}=\text{YOLO}(\mathbf{X}_{I})=[x_{c},y_{c},w,h]^{T},$ (3)
which consists of the detected object’s center coordinates ($x$-axis, $y$-axis), width, and height within the RGB image. Since the optimal beam selection is highly dependent on the direction and position of the transmission target, we use a sequence of bounding box vectors as the extracted visual feature. The objective is to predict the optimal beam index for the next $H$ steps based on the historical $T$ step bounding box vectors, denoted as $\mathbf{B}=[\mathbf{b}[t-T+1],\cdots,\mathbf{b}[t-1]]\in\mathbb{R}^{4\times T}$.
@@ III-B The Backbone Module
The inherent potential of LLMs can be utilized to address the beam prediction task. However, a key challenge lies in aligning the visual feature modality with the textual modality to enable LLMs to effectively comprehend the task. Furthermore, fine-tuning LLMs requires extensive datasets, which is often unrealistic in practical scenarios.
@@
Input Embedding: For each row of $\mathbf{B}$, denoted as $\mathbf{B}^{(i)}\in\mathbb{R}^{1\times T}$ for $i=1,2,3,4$, reversible instance normalization (RevIN)[13] is applied individually to normalize the data, ensuring a mean of 00 and a variance of $1$. RevIN dynamically adjusts the normalization parameters to accommodate variations in the data distribution. Subsequently, $\mathbf{B}^{(i)}$ is segmented into several contiguous overlapping or non-overlapping patches, each of length $L_{p}$. The total number of input patches is given by $\lfloor\frac{T-L_{p}}{S}\rfloor-2$, where $S$ represents the horizontal sliding size. This operation is inspired by techniques in CV, wherein local temporal information is aggregated within each patch to better preserve local semantic features. Finally, a simple linear layer is employed to embed $\mathbf{B}^{(i)}_{P}\in\mathbb{R}^{P\times L_{p}}$ into $\hat{\mathbf{B}}^{(i)}_{P}\in\mathbb{R}^{P\times d_{m}}$.
Patch Reprogramming: Since natural language and input features belong to different modalities, with different ways of representing semantics, LLMs cannot directly process $\hat{\mathbf{B}}^{(i)}_{P}$. To address this, the reprogramming layer maps the input time sequence of visual features into an NLP task, enabling the utilization of LLMs’ reasoning and inference capabilities. A common technique for aligning different modalities is cross-attention[14], which enables interactions between word embeddings and input features by dynamically attending to relevant information across modalities. In this framework, the temporal input features serve as the query, while the word embeddings act as the key and value. However, given that the backbone model is a general-purpose LLM, the original vocabulary of size $V$ is not entirely relevant to our task. Directly aligning inputs features with all words is impractical, as many words do not carry semantic relevance to the task. Therefore, a simple linear layer is employed to extract text prototypes (semantic prototypes) by projecting pre-trained word embeddings $\mathbf{E}\in\mathbb{R}^{V\times D}$ onto a small collection of text prototypes $\mathbf{E}^{\prime}\in\mathbb{R}^{V^{\prime}\times D}$, where $V^{\prime}\ll V$. Here, $D$ is the hidden dimension of the backbone model. This projection effectively reduces the number of words from $V$ to $V^{\prime}$, allowing the temporal input features to align only with these prototypes. For each head $k={1,2,\cdots,K}$, we define:
$\mathbf{Q}_{i}^{(k)}$ $=\hat{\mathbf{B}}^{(i)}_{P}\mathbf{W}_{k}^{Q},$ (4)
$\mathbf{K}_{i}^{(k)}$ $=\mathbf{E}^{\prime}\mathbf{W}_{k}^{K},$ (5)
$\mathbf{V}_{i}^{(k)}$ $=\mathbf{E}^{\prime}\mathbf{W}_{k}^{V},$ (6)
where $\mathbf{W}^{Q}_{k}\in\mathbb{R}^{d_{m}\times\lfloor\frac{d_{m}}{K}\rfloor}$ and $\mathbf{W}^{K}_{k},\mathbf{W}^{V}_{k}\in\mathbb{R}^{D\times\lfloor\frac{d_{m}}{K}\rfloor}$.
The following process adaptively obtains the text descriptions corresponding to patches through a multi-head self-attention mechanism:
@@
$=\mathrm{SOFTMAX}\left(\frac{\mathbf{Q}_{k}^{(i)}{\mathbf{K}_{k}^{(i)}}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{k}^{(i)}.$ (7)
By aggregating each $\mathbf{Z}^{(i)}_{k}\in\mathbb{R}^{P\times d}$ across all heads, we obtain $\mathbf{Z}^{(i)}\in\mathbb{R}^{P\times d_{m}}$. This is then linearly projected to align the hidden dimension with the backbone model, resulting in $\mathbf{O}^{(i)}\in\mathbb{R}^{P\times D}$.
@@
Output Projection: By packaging and forwarding the prompts along with the patch embeddings $\mathbf{O}^{(i)}$ through the frozen LLM, we discard the prefix portion and obtain the output representations. These representations are then flattened and linearly projected to produce the final outputs, $\hat{\mathbf{P}}=[\hat{\mathbf{p}}[t],\cdots,\hat{\mathbf{p}}[t+H-1]]\in\mathbb{R}^{M\times H}$. The index of the dimension corresponding to the maximum value of each $\hat{\mathbf{p}}[t]$, is predicted as the optimal future beam index, given by:
$\hat{m}^{*}[t]=\arg\max_{m\in[1,|\mathbf{\mathcal{F}}|]}\hat{\mathbf{p}}[t].$ (8)

### v2

@@ Abstract
In this paper, we propose BeamLLM, a vision-empowered millimeter-wave (mmWave) beam prediction framework leveraging large language models (LLMs) to enhance the accuracy and robustness of beam prediction. By integrating computer vision (CV) with LLMs’ cross-modal reasoning capabilities, the framework extracts user equipment (UE) positional features from RGB images and aligns visual-temporal features with LLMs’ semantic space through reprogramming techniques. Evaluated on a realistic vehicle-to-infrastructure (V2222I) scenario, BeamLLM achieves $61.01\%$ top-$1$ accuracy and $97.39\%$ top-$3$ accuracy in standard prediction tasks, outperforming traditional deep learning models. In few-shot prediction scenarios, performance degradation is limited to $12.56\%$ (top-$1$) and $5.55\%$ (top-$3$) from time sample $1$ to $10$, demonstrating superior prediction capability.
@@ I Introduction
Millimeter-wave (mmWave) communication offers abundant spectrum above $26$ GHz, enabling high data rates but suffering from severe path loss. To address this, large antenna arrays employ narrow beams to enhance link quality. Effective beamforming requires accurate beam alignment between transmitters and receivers, traditionally achieved via beam training using predefined codebooks. However, this process incurs high overhead, especially in high-mobility scenarios such as V2222X and UAVs. To address this issue, a promising approach is to proactively predict beams by leveraging user-side positioning data and base station (BS) sensing measurements. For instance, RGB images captured by BS cameras provide both spatial location of the UE and rich environmental cues. Predicting beams from historical image sequences involves capturing the complex spatiotemporal interactions within sensing data-making. This task well-suited for DL, which can uncover hidden dynamics and learn mappings between visual input and future beam directions [1, 2, 3, 4]. As a key technique for integrated sensing and communication (ISAC) in $6$G, this approach shows strong potential for efficient mmWave MIMO operation.
@@
Recent breakthroughs in large language models (LLMs), such as GPT-$4$ [5] and DeepSeek [6], have demonstrated remarkable capabilities in contextual reasoning and few-shot generalization abilities. While LLMs are originally designed for natural language processing (NLP), LLMs have shown strong cross-modal learning capabilities, thus extending their applications to other tasks. Inspired by these advantages of the LLMs, several works applying LLMs have been proposed for channel prediction[7], beam prediction [8], and port prediction for fluid antennas[9].
@@
Built on these developments, in this paper, we propose a vision-empowered beam prediction framework, named BeamLLM, which utilizes LLMs to process historical RGB image series, thereby enabling more efficient and adaptive beam prediction. Unlike [8], our method does not rely on historical beam indices or angle of departure (AoD) information. Instead, BeamLLM relies solely on visual features for beam prediction. Furthermore, to ensure robust performance and practical applicability, we validate our framework using real-world measurement datasets, to demonstrate its potential for deployment in real-world scenarios.
@@
The rest of this paper is organized as follows: Section II provides a system model and problem formulation of the beam prediction task. The proposed BeamLLM framework is presented in Section III. Section IV presents simulation results, including performance comparisons with benchmark methods, along with detailed discussions. Finally, we conclude our work in Section V.
Notations: Bold lowercase letters denote vectors (e.g., $\mathbf{x}$), and bold uppercase letters denote matrices (e.g., $\mathbf{X}$). The superscripts $(\cdot)^{\mathsf{T}}$ and $(\cdot)^{\mathsf{H}}$ represent the transpose and Hermitian (conjugate transpose) operations, respectively. The operator $\mathbb{E}[\cdot]$ denotes the statistical expectation, while $|\cdot|_{2}$ denotes the Euclidean norm of a vector, and ||$\cdot$|| ⋅ | returns the magnitude of a complex number. The indicator function $\mathbbm{1}\{\cdot\}$ equals $1$ if the condition inside the braces is true, and 00 otherwise. Unless otherwise specified, $\mathbb{Z}^{+}$, $\mathbb{R}$, and $\mathbb{C}$ denote the sets of positive integer, real, and complex numbers, respectively.
Figure 1: Illustration of the system model, showing a camera-equipped BS communicating with a moving UE, while RGB images are collected to support communications.
@@ II-A System Model
Fig. 1 illustrates the system model considered for vehicle-to-infrastructure (V2222I) mmWave communication. In this model, the BS deploys a mmWave phased-array receiver with $N$ elements of half-wavelength spacing and an RGB camera. The antenna array enables the BS to perform beamforming, while the camera captures images within its field of view at a certain frame rate for sensing and downstream applications.
@@
We assume that the BS has a predefined beamforming codebook =$\mathcal{F}=\{\mathbf{f}_{1},\cdots,\mathbf{f}_{M}\}$, containing $M$ beams, where $\mathbf{f}_{m}\in\mathbb{C}^{N\times 1},m=1,\cdots,M$ represents the $m$-th beamforming vector. The UE is assumed to have a single antenna. At time step $t$, the user transmits a single symbol $s[t]\in\mathbb{C}$ that satisfies the power constraint $\mathbb{E}[|s[t]|^{2}]=P$, where $P$ represents the transmit power. At the BS, the received signal $y[t]$ can be expressed as:
$y[t]=\mathbf{h}^{\mathsf{H}}[t]\mathbf{f}_{m[t]}s[t]+n[t],$ (1)
where $\mathbf{h}[t]$ is the channel vector, $\mathbf{f}_{m[t]}$ is the $m[t]$-th beamforming vector from the codebook in time step $t$, and $n[t]\sim\mathcal{CN}(0,\sigma^{2})$ is the additive white Gaussian noise (AWGN) with variance $\sigma^{2}$.
@@ II-B Problem Formulation
This paper mainly focuses on the beam prediction problem at the BS. Given the available sensing information up to time $t-1$, the BS attempts to determine the optimal beams for $T_{\text{pred}}\in\mathbb{Z}^{+}$ future time steps, specifically for $t,\cdots,(t+T_{\text{pred}}-1)$. We define the optimal beam at time step $t$ as the beam that provides the highest beamforming gain, given by:
$\mathbf{f}_{m^{*}[t]}=\arg\max_{\mathbf{f}_{m[t]}\in\mathcal{F}}|\mathbf{h}^{\mathsf{H}}[t]\mathbf{f}_{m[t]}|^{2}.$ (2)
When perfect CSI knowledge is unavailable, beam training serves as an alternative method for determining the optimal beam. However, with a narrow beam codebook, the training overhead can be significant, and the likelihood of identifying the optimal beam is often low when the pre-beamforming signal-to-noise ratio (SNR) is poor. Because the optimal beam selection at the transmitter and receiver depends on the surrounding environment of the transceiver, our work aims to leverage visual information at the BS to assist beam selection and develop a beam prediction framework. Cameras deliver detailed visual data without requiring wireless resources or feedback signals, unlike radar or wireless-based sensing. As mature, cost-effective sensors, they are easier to implement and deploy. Additionally, cameras can leverage advanced computer vision algorithms for enhanced functionality [1].
@@ III Large Laguage Model-Based Beam Prediction
In this section, we introduce BeamLLM to tackle the vision-assisted beam prediction task outlined in Section II. Fig. 2 illustrates the proposed BeamLLM. The architecture mainly comprises two components, i.e., the visual data feature extraction module and the backbone module.
@@ III-A Visual Data Feature Extraction Module
To process raw RGB data for the vision-aided beam prediction task, we employ the YOLOv4 object detector[10]. We select YOLOv4 due to its proven stability, computational efficiency, and robust performance in real-time object detection tasks, which are critical for practical mmWave beam prediction scenarios. This detector identifies potential UEs within RGB images and extracts bounding box vectors $\mathbf{b}$. For a single image $\mathbf{X}_{I}\in\mathbb{R}^{W_{\text{I}}\times H_{\text{I}}\times C_{\text{I}}}$, where $W_{I}$, $H_{I}$ and $C_{I}$ represent the spatial width, height, and RGB/IR spectral channels respectively, the bounding box vector is given as:
$\mathbf{b}=\text{YOLO}(\mathbf{X}_{I})=[x_{c},y_{c},w,h]^{\mathsf{T}},$ (3)
which consists of the detected object’s center coordinates ($x$-axis, $y$-axis), width, and height within the RGB image. Since the optimal beam selection is highly dependent on the direction and position of the transmission target, we use a sequence of bounding box vectors as the extracted visual feature. The objective is to predict the optimal beam index for the next $T_{\text{pred}}$ steps based on the historical $T_{\text{hist}}$ step bounding box vectors, denoted as $\mathbf{B}=[\mathbf{b}[t-T_{\text{hist}}],\cdots,\mathbf{b}[t-1]]\in\mathbb{R}^{4\times T_{\text{hist}}}$.
@@ III-B The Backbone Module
The inherent potential of LLMs can be utilized to address the beam prediction task. However, a key challenge lies in aligning the visual feature modality with the textual modality to enable LLMs to effectively comprehend the task. Furthermore, training LLMs requires extensive datasets, which is often unrealistic in practical scenarios.
@@
Input Embedding: For each row of $\mathbf{B}$, denoted as $\mathbf{B}^{(i)}\in\mathbb{R}^{1\times T_{\text{hist}}}$ for $i=1,2,3,4$, reversible instance normalization (RevIN)[12] is applied individually to normalize the data, ensuring a mean of 00 and a variance of $1$. RevIN dynamically adjusts the normalization parameters to accommodate variations in the data distribution. Subsequently, $\mathbf{B}^{(i)}$ is segmented into several contiguous overlapping or non-overlapping patches, each of length $L_{p}$. The total number of input patches is given by $\lfloor\frac{T_{\text{hist}}-L_{p}}{S}\rfloor+2$, where $S$ represents the horizontal sliding size. This operation is inspired by techniques in CV, wherein local temporal information is aggregated within each patch to better preserve local semantic features. Finally, a simple linear layer is employed to embed $\mathbf{B}^{(i)}_{P}\in\mathbb{R}^{P\times L_{p}}$ into $\hat{\mathbf{B}}^{(i)}_{P}\in\mathbb{R}^{P\times d_{m}}$.
Patch Reprogramming: To enable LLMs to process visual features $\hat{\mathbf{B}}^{(i)}_{P}$, which differ from natural language in modality, a reprogramming layer maps temporal input features to an NLP task using cross-attention [13]. A linear layer projects pre-trained word embeddings $\mathbf{E}\in\mathbb{R}^{V\times D}$ onto a smaller set of text prototypes $\mathbf{E}^{\prime}\in\mathbb{R}^{V^{\prime}\times D}$, where $V$ and $V^{\prime}$ denote the vocabulary sizes of $\mathbf{E}$ and $\mathbf{E}^{\prime}$, respectively $(V^{\prime}\ll V)$, and $D$ is the hidden dimension. For each attention head $k=1,\ldots,K$, we compute:
$\mathbf{Q}_{k}^{(i)}$ $=\hat{\mathbf{B}}^{(i)}_{P}\mathbf{W}_{k}^{Q},$ (4)
$\mathbf{K}_{k}^{(i)}$ $=\mathbf{E}^{\prime}\mathbf{W}_{k}^{K},$ (5)
$\mathbf{V}_{k}^{(i)}$ $=\mathbf{E}^{\prime}\mathbf{W}_{k}^{V},$ (6)
where $\mathbf{W}^{Q}_{k}\in\mathbb{R}^{d_{m}\times\lfloor\frac{d_{m}}{K}\rfloor}$ and $\mathbf{W}^{K}_{k},\mathbf{W}^{V}_{k}\in\mathbb{R}^{D\times\lfloor\frac{d_{m}}{K}\rfloor}$. The following process adaptively obtains the text descriptions corresponding to patches through a multi-head self-attention mechanism:
@@
$=\mathrm{SOFTMAX}\left(\frac{\mathbf{Q}_{k}^{(i)}{\mathbf{K}_{k}^{(i)}}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{k}^{(i)},$ (7)
where $\mathrm{ATTENTION}(\cdot)$ represents the scaled dot-product attention mechanism and $\mathrm{SOFTMAX}(\cdot)$ normalizes attention scores into a probability distribution, respectively. By aggregating each $\mathbf{Z}^{(i)}_{k}\in\mathbb{R}^{P\times d}$ across all heads, we obtain $\mathbf{Z}^{(i)}\in\mathbb{R}^{P\times d_{m}}$. This is then linearly projected to align the hidden dimension with the backbone model, resulting in $\mathbf{O}^{(i)}\in\mathbb{R}^{P\times D}$.
@@
Output Projection: By packaging and forwarding the prompts along with the patch embeddings $\mathbf{O}^{(i)}$ through the frozen LLM, we discard the prefix portion and obtain the output representations. These representations are then flattened and linearly projected to produce the final outputs, $\hat{\mathbf{P}}=[\hat{\mathbf{p}}[t],\cdots,\hat{\mathbf{p}}[t+T_{\text{pred}}-1]]\in\mathbb{R}^{M\times T_{\text{pred}}}$. The index of the dimension corresponding to the maximum value of each $\hat{\mathbf{p}}[t]$, is predicted as the optimal future beam index, given by:
$\hat{m}^{*}[t]=\arg\max_{m}\hat{\mathbf{p}}[t].$ (8)